OpenStudy (anonymous):

okay I part of this question but can you help me walk though it again? mutifply and simplfly (see drawing)

OpenStudy (anonymous):

\[\sqrt{3}(4+\sqrt{12})\]

OpenStudy (anonymous):

\(\sqrt{3}\times 4=4\sqrt{3}\)

OpenStudy (anonymous):

\(\sqrt{3}\times \sqrt{12}=\sqrt{36}=6\)

OpenStudy (anonymous):

and so final answer is \(4\sqrt3 + 6\)

OpenStudy (anonymous):

thank you, for some reason the equation thing is not working right but thank you!